### v1

# Early warning signals of the tipping point in strongly interacting Rydberg atoms

22footnotetext: J.Z, Z.K.L, L.H.Z contribute equally to this work.11footnotetext: dds@ustc.edu.cn
Early warning signals of the tipping point in strongly interacting Rydberg atoms
Jun Zhang1,2,†    Zong-Kai Liu1,2,†    Li-Hua Zhang1,2,†    Bang Liu1,2    Zheng-Yuan Zhang1,2    Shi-Yao Shao1,2    Qing Li1,2    Han-Chao Chen1,2    Yu Ma1,2    Tian-Yu Han1,2    Qi-Feng Wang1,2    C. Stuart Adams3    Dong-Sheng Ding1,2,⋆    Bao-Sen Shi1,2 1Key Laboratory of Quantum Information, University of Science and Technology of China, Hefei, Anhui 230026, China. 2Synergetic Innovation Center of Quantum Information and Quantum Physics, University of Science and Technology of China, Hefei, Anhui 230026, China. 3Department of Physics, Joint Quantum Centre (JQC) Durham-Newcastle,
Durham University, South Road, Durham DH1 3LE, United Kingdom.
(May 1, 2024)
The identification of tipping points is essential for prediction of collapses or other sudden changes in complex systems. Applications include studies of ecology, thermodynamics, climatology, and epidemiology. However, detecting early signs of proximity to a tipping is made challenging by complexity and non-linearity. Strongly interacting Rydberg atom gases offer a model systems that offer both complexity and non-linearity, including phase transition and critical slowing down. Here, via an external probe we observe prior warning of the proximity of a phase transition of Rydberg thermal gases. This warning signal is manifested as a cessation of the variance growth with increasing probe intensity. We also observed the dynamics of the critical slowing down behavior versus different time scales, driving intensities, and atomic densities, thus providing insights into the study of a Rydberg atom system’s critical behavior. Our experiment suggests that the full critical slowing down dynamics of strongly-interacting Rydberg atoms can be probed systematically, thus providing a benchmark with which to identify critical phenomena in quantum many-body systems.
As a result of their inherent interconnectedness and nonlinear dynamics, complex systems have tipping points at which small changes in a specific component or parameter can trigger a disproportionately large response in the entire system [1, 2]. Understanding and identifying these tipping points are essential steps in enabling complex systems to be managed effectively [3, 4, 5, 6, 7]. Effective management could allow us to anticipate and respond to potential tipping points and to implement strategies that can avoid or mitigate their detrimental effects [3, 8, 9]. Furthermore, complex systems often show self-organization [10, 11] and emergent properties [12], which means that their behavior at the collective level cannot be predicted solely based on an understanding of the individual components. As the interactions and dependencies between the components become stronger near the tipping point, the system will then become more susceptible to abrupt shifts [13, 14, 15]. Various factors can contribute to the existence of tipping points in different complex systems; examples include epileptic seizures in medicine [16, 17], systemic market crashes in financial networks [18, 19], climate in nature [20, 21, 22], and phase transitions in physics [23, 24, 25]. Several significant works in the literature have reported on building of theoretical models to predict tipping points [4, 26, 27], with examples that include landscape-flux theory [28, 29], deep learning methods [30, 31], and the critical slowing down method [21, 32, 22], in which early warning signals are introduced.
The long-range interactions that occur between Rydberg atoms provide us with the ability to simulate large scale systems and study non-equilibrium phase transitions, self-organized criticality, and other phenomena [33, 34, 10, 35, 36, 37]. Rydberg atoms capture the interconnectedness between the different components in a system and also provide strong nonlinearity, which is particularly relevant for complex systems [38, 39, 40, 41, 42]. Therefore, the study of Rydberg atom systems aids in understanding of emergent behaviors and also provides insights into the dynamics of complex systems [43, 44, 45, 46, 47, 48]. However, the complex interconnectedness and nonlinearity, the nonstationary dynamics, and the uncertainty evolution of such systems mean that they are difficult to control and manipulate [49, 50, 51, 52]. For example, even small changes in the system parameters can lead to significant responses within the entire atomic system, where the behavior of one component can have far-reaching effects on the overall system dynamics, including inducing rich emergent properties and self-organization dynamics [10]. This also represents a widespread problem among other complex systems. However, no experiments on prediction of the tipping points in Rydberg atom systems have been reported to date, and these systems are worthy of study.
Figure 1: Physical model and early warning signals of the tipping point. (a) A phase transition in a complex network, in which phase 1 and phase 2 represent the different states of matter below and above the tipping point, respectively. These two phases show different properties, and we use the interconnectedness link here between the particles in phase 2, as compared with no available link in phase 1. (b) Two-level atom model including the ground state $\left|g\right\rangle$ and the Rydberg state $\left|r_{1}\right\rangle$, along with the spontaneous radiation rate $\Gamma_{1}$. The atoms are driven from the ground state to the Rydberg state by a laser with a Rabi frequency $\Omega_{1}$ and detuning $\Delta_{1}$. The atoms are then driven by a modulated microwave (MW) electric field with a Rabi frequency of $\Omega_{2}$. The Rydberg energy level is then modified by the many-body interaction strength $V={C_{6}}/{R^{6}}$ [where $C_{6}$ is the van der Waals coefficient and $R$ is the distance between the Rydberg atoms]. (c) Phase transition of the Rydberg atom system. The Rydberg atom population $\rho_{r_{1}r_{1}}$ shows a sudden change versus the square of the Rabi frequency ${\Omega_{1}}^{2}$, where ${\Omega}^{2}_{1,c}$ corresponds to the tipping point. (d) Measured transmission of the probe beam versus the parameter ${\Omega_{p}}^{2}$ under the MW field driving condition. (e) Measured variance from (d). In this process, there is an early warning signal regime, which is marked as the blue shaded area. The variance in this regime is different to the corresponding variance far away from the tipping point, manifesting as a nonlinear growth property. The black dotted line is a linear curve, which is included as a guide for the eye. We fitted the trend of the variance using a solid red line, where the fit function is $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$ ($p$ = -4.44$\times$10-7 $\pm$ -4.16$\times$10-10, $q$ = -3.89$\times$10-9 $\pm$ 3.72$\times$10-10, $s$ = 2.01$\times$10-6 $\pm$ 1.41$\times$10-9, $\alpha$ = 8.50 $\pm$ 0.04).
Here, we report an experiment in which the tipping point of interacting Rydberg atoms can be predicted by driving the system using a resonant microwave (MW) electric field. We observe that early warning signals are manifested as a symptom regime in which the variance in the transmission no longer grows linearly. By using a nondestructive measurement method with the electromagnetically induced transparency (EIT) detection method, we can coherently map the status of the Rydberg atoms with respect to the probe field. The two metastable states of the non-equilibrium system will display different responses because the non-interacting Rydberg atoms demonstrate relatively small dissipation when compared with that of the interacting Rydberg atoms. Interactions make the Rydberg atoms difficult to synchronize under the driving MW electric field. Therefore, in the experiments, we can identify the point at which the system undergoes a phase transition directly. This work provides predictive capabilities with regard to the behavior of complex systems and will enable better early responses to potential tipping points.
Results
Warning signal model
The inherent interconnectedness and the nonlinear interactions within complex systems make it difficult to predict sudden changes in these systems. Modeling of complex systems can help us to find early warning signals before these sudden changes, as illustrated in Fig. 1(a). The Rydberg atom system represents a good platform for study of complex systems because of the strong interactions that occur between the Rydberg atoms, and there is a non-equilibrium phase transition for Rydberg atoms from the non-interacting phase (NI phase) to the interacting phase (I phase), as illustrated in Fig. 1(b,c).
The Rydberg atom system consists of $N$ interacting three-level atoms with a ground state $\left|g\right\rangle$, a Rydberg state $\left|r_{1}\right\rangle$ with a decay rate of $\Gamma_{1}$, and a Rydberg state $\left|r_{2}\right\rangle$ with a decay rate of $\Gamma_{2}$, as depicted in Fig. 1(b); further details about the experiment can be found in the Methods section. A two-photon process excites the atom from the ground state $\left|g\right\rangle$ to the Rydberg state $\left|r_{1}\right\rangle$ with a Rabi frequency of $\Omega_{1}$ and detuning of $\Delta_{1}$. The intermediate energy level of this two-photon process is eliminated adiabatically and is not shown in Fig. 1(b).
A dual-tone MW field drives the atoms between the Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and this MW field takes the form of $E\rm_{MW}$$\cos(\omega_{1}t)+E\rm_{MW}\cos{[(\omega_{1}+\delta f)t]}$, where $E\rm_{MW}$ represents the amplitude, $\omega_{1}$ is the resonant frequency with respect to the transition between the two Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and $\delta f$ is the MW frequency detuning from resonance. After application of the rotation wave approximation, this dual-tone MW field turns into a periodic driving field with frequency $\delta f$. When the MW field is off and the intensity profile of the probe laser is scanned, a non-equilibrium phase transition occurs that is characterized by a sudden increase in the population of the Rydberg atoms $\rho_{r1r1}$, as shown in Fig. 1(c). When the dual-tone MW field is on, an oscillation effect occurs in the transmission spectrum. The MW field serves as a perturbation in this case and the spectrum represents the response of the Rydberg atoms to this perturbation, as illustrated in Fig. 1(d). When the Rabi frequency of the probe field is low and the system is outside the warning regime, which is shaded in blue in Fig. 1(d), the Rydberg atom system is then a linear system and its response to the MW field is also linear, as indicated by the linear growth observed in the variance as the square of the Rabi frequency $\Omega_{p}^{2}$ increases; see the Methods section for the definition of the variance. However, near the sharp jump in the transmission, there is a warning regime region colored blue, where the variance in the transmission decreases as $\Omega_{p}^{2}$ increases, as shown in Fig. 1(e). This variance, which deviates from the previously linear increase, represents a warning signal.
A toy numerical model with an interaction-dependent energy shift and energy broadening can reveal the physical mechanism of our experiment.
The master equation is
$\partial_{t}\hat{\rho}=i[\hat{H},\hat{\rho}]+\mathcal{L}_{r_{1}}[\hat{\rho}]+\mathcal{L}_{r_{2}}[\hat{\rho}]$ (1)
where $\hat{H}$ is the Hamiltonian and $\mathcal{L}_{\{r_{1},r_{2}\}}$ is the Lindblad jump operator. The Hamiltonian of the system after adiabatic elimination of the intermediate energy level between $\left|g\right\rangle$ and $\left|r\right\rangle$ is given by
$\hat{H}$ $=\frac{1}{2}\sum_{i}\left(\Omega_{1}\sigma_{i}^{gr_{1}}+\Omega_{2}\sigma_{i}^{r_{1}r_{2}}+h.c.\right)+\sum_{i}\left(\Delta_{1}n_{i}^{r_{1}}+\Delta_{2}n_{i}^{r_{2}}\right)$
$+\frac{1}{2}\sum_{i\neq j}(V_{ij}^{r_{1}r_{2}}n_{i}^{r_{1}}n_{j}^{r_{2}}+V_{ij}^{r_{2}r_{1}}n_{i}^{r_{2}}n_{j}^{r_{1}}+V_{ij}^{r_{1}r_{1}}n_{i}^{r_{1}}n_{j}^{r_{1}}$
$+V_{ij}^{r_{2}r_{2}}n_{i}^{r_{2}}n_{j}^{r_{2}})$
where $\sigma_{i}^{gr}$ ($r={r_{1},r_{2}}$) represents the $i$-th atom transition between the ground state $\left|g\right\rangle$ and the Rydberg state $\left|r\right\rangle$; $n_{i}^{r_{1},r_{2}}$ are the population operators of the two Rydberg energy levels $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$; $V_{ij}^{r_{1}r_{2}}$ and $V_{ij}^{r_{2}r_{1}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1,2}\right\rangle$ and $\left|r_{2,1}\right\rangle$, respectively; and $V_{ij}^{r_{1}r_{1}}$ and $V_{ij}^{r_{2}r_{2}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, respectively.
The Lindblad jump terms are:
$\mathcal{L}_{r}=(\gamma_{r}/2)\sum_{i}(\hat{\sigma}_{i}^{rg}\hat{\rho}\hat{\sigma}^{gr}_{i}-\{\hat{n}_{i}^{r},\hat{\rho}\}),$ (2)
where these terms represent the decay process from the Rydberg state $\left|r\right\rangle$ to the ground state $\left|g\right\rangle$.
Note here that the mean-field approximation is used and that the interaction terms ($V_{ij}^{r_{1}r_{2}}$, $V_{ij}^{r_{2}r_{1}}$, $V_{ij}^{r_{1}r_{1}}$, and $V_{ij}^{r_{2}r_{2}}$) are replaced by a single atom immersed in a field formed by these interactions between other atoms and this atom. Therefore, the task of solving the dynamic many-body problem is reduced to solving the dynamic problem for one single atom in that field. Besides, before the non-equilibrium phase transition occurs, the interaction between Rydberg atoms dissipates the atoms, which makes the dissipation ahead of the phase transition (the deviation from linear growth of the variance before the tipping point as shown in Fig. 1(e).). To model this process, we introduce a density-dependent dissipation in the mean-field approximation, and the detuning and the decay are replaced such that $\Delta_{1}\rightarrow\Delta_{1}-V\rho_{r_{1}r_{1}}^{\beta}$ and $\Gamma_{r_{1}}\rightarrow\Gamma_{r_{1}}-V\rho_{r_{1}r_{1}}^{\epsilon}$ [53, 23, 39, 10, 34], respectively, where $\beta$ and $\epsilon$ are exponent factors. By solving the master equation using the mean-field approximation and expanding the solution $\rho_{r_{1}r_{1}}$ as a function of $\Omega_{1}$ to the second order, the transmission against the intensity of the probe field $(\Omega_{1}/2\pi)^{2}$ is obtained. Here, $\beta=1$ and $\epsilon=3$ to enable better modeling of the results.
Next, we apply the MW field with dual tone frequencies to the theoretical model and simulate the transmission, where the MW field takes the form of $\Omega_{2}=E_{\mathrm{A}}(1+\mathrm{Cos}(\delta{f}t))$, and $E_{\mathrm{A}}$ is the magnitude of the Rabi frequency of the dual tone MW field. The simulated results for $\rho_{r_{1}r_{1}}$ versus $(\Omega_{1}/2\pi)^{2}$ are shown in Fig. 2(a-b), where panel (a) corresponds to the case of the phase transition without the MW field ($E_{\mathrm{A}}=0$) and panel (b) shows the case with the applied MW field ($E_{\mathrm{A}}\neq 0$). Figure 2(c) shows the variance that was calculated from Fig. 2(b), with reduced behavior in the region where 2.2 ($2.2\text{ }(\mathrm{MHz})^{2}<(\Omega_{1}/2\pi)^{2}<3.5\text{ }(\mathrm{MHz})^{2}$.
Figure 2: Theoretical and experimental results without and with the driving MW field. (a), (b) Theoretical simulations of the population $\rho_{r_{1}r_{1}}$ of the Rydberg atoms with different electric field amplitudes: (a) $E_{\mathrm{A}}$ = 0 $\mathrm{mV/cm}$ and (b) $E_{\mathrm{A}}\neq$ 0 $\mathrm{mV/cm}$. (c) Variance calculated from (b). (d), (e) Experimental transmission spectra under the different MW field amplitudes: (d) $E\rm_{MW}$ = 0 $\mathrm{mV/cm}$ and (e) $E\rm_{MW}$ = 1.64 $\mathrm{mV/cm}$. In this case, the scanning period Ts = 1 ms. (f) Measured variance from (e).
Transmission spectra under the dual-tone MW field driving conditions
By scanning the probe field intensity, we can observe the non-equilibrium phase transition in the Rydberg atom system, as illustrated in Fig. 2(d). To subtract the early warning signals before the tipping point is reached, we experimentally apply a dual-tone MW electric field to drive the transition between the two different Rydberg states $|50D_{5/2}\rangle$ and $|51P_{3/2}\rangle$, with a center frequency of $2\pi\times 17.041$ GHz and a frequency difference of $\delta f=2\pi\times 0.2$ MHz. The MW field induces a small ac Stark shift that moves the tipping point and varies the Rydberg atom population. The MW field causes the tipping point to move forward when compared with the case without the MW field, and the oscillation appears in the transmission spectrum in Fig. 2(e). By calculating the variance from the oscillated transmission, we observe a reduced variance when $(\Omega_{p}/2\pi)^{2}>26\text{ }(\mathrm{MHz})^{2}$, indicating the existence of an early warning signal, as shown in Fig. 2(f). The experimental results show good agreement with the expectations of the theoretical simulations.
From the results shown in Fig. 2(f), the measured variance is reduced nonlinearly when $(\Omega_{p}/2\pi)^{2}>27.5\text{ }(\mathrm{MHz})^{2}$. The increased $\Omega_{p}$ adds a number of excited Rydberg atoms, which makes the system dimension higher and then induces a strong critical slowing down effect. This occurs because, in a relatively low-dimensional system where the number of degrees of freedom is limited, localized interactions between neighboring Rydberg atoms are dominant. As a result, the critical slowing down process in a low-dimensional system is primarily influenced by these localized interactions. In this case, all the excited Rydberg atoms in the system can be synchronized rapidly, thus allowing faster recovery from the perturbation. In contrast, a high-dimensional system has large numbers of interacting Rydberg atoms, which results in a rich interconnectedness within the system. In this system, the critical slowing down effects are amplified because of the extended reach and the influence of the long-range interactions. Perturbations take longer times to propagate through the system, and the process of recovery from disturbances thus becomes slower.
Figure 3: Time evolution of the system during the phase transition process. (a)-(c) Double well model used to describe the phase transition and the critical slowing down phenomenon. With increasing ${\Omega_{\mathrm{p}}}^{2}$, the system is initially in phase 1 (${\Omega_{\mathrm{p}}}^{2}<{\Omega_{\mathrm{p,c}}}^{2}$, where ${\Omega_{\mathrm{p,c}}}^{2}$ is the tipping point), and it then evolves from state A to state B when close to the tipping point, before finally making the transition to phase 2 ((${\Omega_{\mathrm{p}}}^{2}>{\Omega_{\mathrm{p,c}}}^{2}$), state C). In this process, the system states A, B, and C have different interaction-induced dissipations $\gamma_{A}$, $\gamma_{B}$, and $\gamma_{C}$, respectively, where $\gamma_{C}>\gamma_{B}>\gamma_{A}$. (d)-(f) Measured transmission spectra and their corresponding variances within scanning times of (d) 0.25 ms, (e) 0.5 ms, and (f) 2 ms.
Transmission spectra versus scanning time
The non-equilibrium phase transition in the Rydberg atom system can be modeled using a double-well transition, with each well standing for one bi-stable state, as depicted in Fig. 3(a)-(c). In the left well, phase 1 represents the few-body state, while in the right well, phase 2 corresponds to the many-body state. The dissipation of phase 2 is greater than that of phase 1 because of the interactions among the Rydberg atoms. When the population of Rydberg atoms is increased by $\Omega_{p}^{2}$, a phase transition occurs from the left well to the right well that crosses the barrier of the double well, as illustrated in Fig. 3(b). Near the tipping point $\Omega_{p,c}^{2}$, the critical slowing down phenomenon occurs; this term refers to an effect where the system’s response time to the driving MW field becomes slower (here, the MW field is regarded as a perturbation to the system). After the phase transition, the system is then in the I-phase, and the strong interactions between the Rydberg atoms make the system’s dissipation so large that the variance drops to a very low level.
To observe the slower dynamics that occur near the tipping point, we changed the scanning time (Ts) of the triangular wave loaded onto the acousto-optic modulator (AOM) and recorded the dynamics within the warning regime. As shown in Fig. 3(d)-(f), we found that the warning regime gradually becomes apparent as the scanning rate is gradually reduced.
A fast scan accumulates a small number of Rydberg atoms because of the small interval time. This causes the system to have low dissipation until the phase transition occurs, which means that the warning regime is almost invisible [see Fig. 3(d) and Fig. 3(e) for the scanning times of Ts= 0.25 ms and Ts= 0.5 ms, respectively]. As the light intensity increases, the system’s response to the external perturbation follows the increasing light intensity well.
Figure 4: Trend of variance at different MW field intensities. (a) Variance measured at the different MW field intensities of $E_{\mathrm{MW}}$=1.80 mV/cm (blue), $E_{\mathrm{MW}}$=1.15 mV/cm (light green), and $E_{\mathrm{MW}}$=0.66 mV/cm (brown); see the Methods section for further details. To provide a better comparison of the signals at the different microwave intensities, we normalized the variance of the transmission for easier reading. The variance at 0.66 mV/cm is amplified by a factor of four and the variance at 1.15 mV/cm is amplified by a factor of 1.5, but the variance at 1.80 mV/cm is not amplified. (b) Theoretical simulations at the different modulation strengths of $E_{\mathrm{A}}$ = 2 (blue), $E_{\mathrm{A}}$ = 0.4 (black), and $E_{\mathrm{A}}$ = 0.1 (brown), where we normalized the simulated results using an amplification factor of 10 at $E_{\mathrm{A}}$ = 0.4 and an amplification factor of 20 at $E_{\mathrm{A}}$ = 0.1; the variance at $E_{\mathrm{A}}$ = 2 was not amplified. The red points correspond to the points where the variance diverged from the linear function in each case. In these processes, the tipping points for the different MW field intensities are calibrated to be in the same position for ease of comparison by slightly adjusting the coupling detuning.
As the scanning time increases, the system can accumulate increasing numbers of Rydberg atoms for a relatively large time interval, and the dissipation of these interacting Rydberg atoms competes with the coherent driving of the laser and the MW field. The dissipation becomes dominant when the system approaches the tipping point, as indicated by the transition process $A\rightarrow B\rightarrow C$ shown in Fig. 3(b). This causes the variance to diverge from the linear increasing trend, and thus the warning regime gradually becomes visible [see Fig. 3(f) for the scanning time of Ts = 2 ms].
Transmission spectra versus the MW field intensity
To study the early warning signals under various MW field intensities, we alter the MW field intensity and record the corresponding variance; see the results in Fig. 4(a). When $E\rm_{MW}$ = 0.66 mV/cm, the variance is almost linear versus the probe field intensity. During this process, the MW field intensity reflects the magnitude of the perturbation. For a small perturbation, the Rydberg atoms can follow the driving MW field even though the system is very close to the tipping point, which means that the warning regime is unclear. When we increase the MW field amplitude, the variance diverges from the linear curve, as shown in the cases where $E\rm_{MW}$ = 1.15 mV/cm and $E\rm_{MW}$ = 1.80 mV/cm. The system then cannot follow the driving action of the MW field, which makes the warning regime become more apparent.
These results can be explained as the process by which the system is synchronized or not when driven by the MW field. The non-interacting Rydberg atoms can cooperate well and are further synchronized with the MW field, while the interacting Rydberg atoms in the system are difficult to synchronize because of the effect of the critical slowing down process. We also simulated the results with different MW field intensities, with results as shown in Fig. 4(b). The simulated results show the same trend as the experimental measurements.
Figure 5: Variance measured at different temperatures. We measured the different variance characteristics at temperatures of (a) T=47.2C, (b) T=47.7C, (c) T=48.2C, and (d) T=49.1C. The results show that as the temperature increases, the divergence of the variance from the linear regime becomes more obvious. The early warning signal regimes shown in (a) and (d) are marked as the blue shaded areas for comparison. The solid red lines are the fittings with the function $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$, where $\alpha$ = 16.46 $\pm$ 0.22 for T=47.2C, $\alpha$ = 12.69$\pm$ 0.10 for T=47.7C, $\alpha$ = 10.89 $\pm$ 0.07 for T=48.2C, and $\alpha$ = 8.59 $\pm$ 0.07 for T=49.2C.
Transmission spectra at different temperatures
In our experiments, we also studied how the interaction between the Rydberg atoms affects the warning signals. A higher temperature leads to a higher interaction level, and vice versa. We changed the atomic density by varying the temperature and measured the variance of the oscillated transmission. Figure 5(a)-(d) show the measured results. We found that the variance before the phase transition was decreasing gradually as the temperature increased. For comparison, we found that the early warning signal regime shown in Fig. 5(d) was broader than that shown in Fig. 5(a), as given by ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=3.0\text{ }(\mathrm{MHz})^{2}$ at T = 47.2C and ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=1.5\text{ }(\mathrm{MHz})^{2}$ at T = 49.1C. Therefore, it is easier to identify the early warning signal at higher temperatures. This is explained by the fact that the interaction-induced dissipation causes the system to be difficult to synchronize when driven by the MW field, and it then leaves the linear response regime. This enhanced critical slowing down process in high-dimensional systems can be attributed to the increased system complexity and the interdependences among the Rydberg atoms.
The early warning signals of the Rydberg atom systems under different spatial structure conditions and scales [from their distinct densities] can also be probed, and the measured exponent $\alpha$ varies with the temperature [e.g., for T = 47.2C, $\alpha$ = 16.46 $\pm$ 0.22 in Fig. 5(a), but $\alpha$ = 8.59 $\pm$ 0.07 for T = 49.2C in Fig. 5(d)], thus indicating distinct critical dynamics before occurrence of a phase transition. This confirms that the early warning signal arises from the interactions between the Rydberg atoms, even when the system is still in the NI phase. This finding is consistent with the fact that the non-equilibrium phase transition in the Rydberg atom system occurs because of the interactions among a few local atoms, which eventually spreads to involve the atoms on a global scale. Using the MW field, we can detect small numbers of locally interacting Rydberg atoms, even without causing a phase transition to occur in the system.
Discussion
It should be noted that the measured variance in our experiment was reduced in the early warning regime, which differs from the increased variance that was reported previously [30, 31, 16]. Here, the variance is calculated from the transmission that occurs under the driving action of the external field, and it is proportional to the amplitude of the oscillation. In this way, the interaction-induced dissipation signal in the Rydberg atom system can be amplified, thus allowing us to characterize the full dynamics of the critical slowing down process. The early warning signals for the phase transition cause us to conclude that the interaction-induced dissipation appears first, and the phase transition occurs after that. This is also consistent with the results from our mean-field method, in which the exponent factor $\epsilon$ was greater than $\beta$.
The critical slowing down process of the Rydberg atoms is a type of nonlinear dynamic behavior that can provide technologies and methods to enable study of other nonlinear systems. The interactions within the Rydberg atom system can help us to understand and study the relationships and the coupling effects among the different factors when studying the early warning signals that occur in other complex systems, including atmospheric systems, ecological systems, and socio-economic systems. Most importantly, the technological method for amplification of the signal of the critical slowing down process of Rydberg atoms can provide a new perspective to enable the development and application of prediction strategies and techniques for use in other complex systems. In addition, the collective behavior and self-organization phenomena among the atoms when driven by the MW field can help us to study the essence and the mechanisms of the self-organization phenomena during intervention by external factors, in fields as diverse as traffic flow, urban planning, and immune systems.
In summary, we have observed the full dynamics of the early warning signals at the tipping point for strongly interacting Rydberg atoms. Because of the distinct dissipation of the metastable states of the Rydberg atom system, the responses before and after the tipping point can be probed successfully using the external MW field. The critical slowing down process causes the response to the MW field to deviate from linear growth before the tipping point, which shows good agreement with the theoretical observations. This experiment has helped us to study the critical slowing down that occurs near the criticality of the Rydberg atom system, and provides a deeper understanding of the critical slowing down behavior in other complex systems. This may also promote practical developments in detection of early warnings before catastrophic events in other complex systems.
Acknowledgements
We acknowledge funding from the National Key R&D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY020200), the major science and technology projects in Anhui Province (Grant No. 202203a13010001)
Author contributions statement
D-S.D. conceived the idea. J.Z implemented the physical experiments. L-H.Z processed the microwave signal. D-S.D. and Z-K.L. conducted the theoretical model. D-S.D., J.Z, and Z-K.L. wrote the manuscript. All authors contributed to the experiment preparation and data discussions regarding the results. D-S.D. support this project.
Competing interests
The authors declare no competing interests.
Data and materials availability
All study data are included in the article.